OpenStudy (anonymous):

Find 30% of 70. (1point) 40 21 2.1 0.4

OpenStudy (anonymous):

of in mathematics means multiply...0.30 * 70 = ?

OpenStudy (anonymous):

I you multiply 7 times 3 you'll get your answer.

OpenStudy (amoodarya):

30% of something = 30/100 =3/10 of something

OpenStudy (anonymous):

No...multiply 7 * 0.3 will give the answer...watch out for those pesky decimals

OpenStudy (amoodarya):

it is sufficient to multiply that number by 3/10

OpenStudy (anonymous):

That too...there are several ways... someone above gave an incorrect method

OpenStudy (anonymous):

:S